# Reinforcing Multimodal Understanding and Generation with Dual Self-rewards

### v2

@@ 4.1 Experimental Setup
We used the training set of T2I-CompBench [15] for the text-to-image generation part, which contains 5,600 text prompts. For the visual understanding part, we randomly sampled 2,800 images from the JourneyDB [37] and COCO118K [38] derived from COCO Caption [39] respectively. It is worth noting that the image and text training data are non-parallel, meaning that we did not use any annotated image-text pairs to train the model.
@@
We employ the unified LMM Janus-Pro-7B [7] as our backbone, optimizing it with our dual self-reward mechanism in a self-supervised manner. We compare our tuned model with state-of-the-art (SOTA) models on visual understanding and generation tasks, including models that are designed only for understanding such as LLaVA [2] and InstructBLIP-7B [40], models designed only for generation tasks like CoMat [41] and LDM [14], as well as unified LMMs that can perform both tasks such as Chameleon [31] and Show-o [5]. All baselines can be found in Table 1, Table 2, and Table 3.
@@
We evaluate our method on visual understanding and generation benchmarks and follow their default settings. For visual generation, we use T2I-CompBench [15] and GenEval [16] as the evaluation benchmarks. We test on their original text prompts without any modification. T2I-CompBench comprises 6,000 compositional text prompts evaluating 3 categories (attribute binding, object relationships, and complex compositions) and 6 sub-categories (color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions). GenEval [16] includes 6 tasks (single object, two objects, counting, colors, position, and color attributes). For visual understanding, we use HallusionBench (denoted as HalluBench) [42], LLaVABench, POPE [43], MMB [44], SEEDBench-IMG (denoted as SEED) [45], and MMMU [46].
Implementation Details
4.2 Results on Visual Generation
We compare our model with the SOTA diffusion models and unified LMMs on the T2I-CompBench and GenEval benchmarks. The results are shown in Table 1 and Table 2, respectively. Our trained model is denoted as “Janus-Pro-7B + DSR”, where the model is optimized with DSR and SimPO in the unified strategy.
The results show that our model achieves substantial improvements over the baseline model, with an average increase of 11.68% on T2I-CompBench and overall 5% on GenEval. On T2I-CompBench, our model remarkably improves the performance in all categories, especially in attribute binding with an average improvement of 20%. In spatial relationships and complex compositions, our model also achieves impressive increases of 5% and 3%, respectively. Our model demonstrates the best performance in shape binding, texture binding, and complex compositions, and achieves the second best performance in color binding. On GenEval, our model improves over the baseline model, especially in counting and color attributes with an increase of 13% and 7%. It outperforms the diffusion models overall, achieving the best performance. These results demonstrate the effectiveness and superiority of our method in improving the text-to-image generation capabilities of LMMs, despite the fact that our model is optimized in a completely self-supervised manner without any external supervision or additional data.
Table 1: Evaluation on T2I-CompBench. Und. and Gen. denote “understanding” and “generation”, respectively. Higher ($\uparrow$) values indicate better performance. The best score is in bold, with the second best score underlined. Line in blue is our model.
@@ 4.2 Results on Visual Generation
StrucDiffusion [47] 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
CompDiffusion [48] 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Attend&Excite [49] 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
PixArt-$\alpha$ [50] 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
CoMat [41] 0.7827 0.5329 0.6468 0.2428 0.3187 0.3680
SD-v1.5 [14] 0.3758 0.3713 0.4186 0.1165 0.3112 0.3047
SD-XL-base-1.0 [51] 0.5879 0.4687 0.5299 0.2131 0.3119 0.3237
FLUX.1 [52] 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703
@@
Show-o [5] 0.56 0.41 0.46 0.20 0.30 0.29
Emu3 [4] 0.7544 0.5706 0.7164
Janus-Pro-7B [7] 0.6426 0.3487 0.4848 0.2061 0.3086 0.3510
@@
Table 2: Evaluation on GenEval.
@@
LlamaGen [3] 0.71 0.34 0.21 0.58 0.07 0.04 0.32
LDM [14] 0.92 0.29 0.23 0.70 0.02 0.05 0.37
SDv1.5 [14] 0.97 0.38 0.35 0.76 0.04 0.06 0.43
PixArt-$\alpha$ [50] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.1 [14] 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 [53] 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu3-Gen [4] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL [51] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 [54] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium [55] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
@@
SEED-X [56] 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Show-o [5] 0.95 0.52 0.49 0.82 0.11 0.28 0.53
D-DiT [57] 0.97 0.80 0.54 0.76 0.32 0.50 0.65
LWM [58] 0.93 0.41 0.46 0.79 0.09 0.15 0.47
Transfusion [59] 0.63
ILLUME [12] 0.99 0.86 0.45 0.71 0.39 0.28 0.61
TokenFlow-XL [58] 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Chameleon [31] 0.39
Janus [6] 0.95 0.62 0.28 0.85 0.45 0.42 0.60
Janus-Pro-1B [7] 0.99 0.82 0.48 0.90 0.62 0.57 0.73
Janus-Pro-7B [7] 0.97 0.88 0.57 0.90 0.77 0.64 0.79
@@
4.3 Results on Visual Understanding
For visual understanding, we test the models on the prevalent benchmarks, including HallusionBench (denoted as HalluBench) [42], LLaVABench, POPE [43], MMB [44], SEEDBench-IMG (denoted as SEED) [45], and MMMU [46], which are widely used to evaluate the understanding capabilities of LMMs. The experimental results can be found in Table 3.
On HallusionBench and LLaVABench, our proposed method achieves substantial improvements over Janus-Pro-7B, the baseline model, with an increase of 1.9% and 7.5%, respectively. On other benchmarks, our model keeps the performance of Janus-Pro-7B, which is already a strong baseline. The results demonstrate that our method can improve the understanding capabilities of LMMs and maintain the original advanced performance, while remarkably improving the generation capabilities.
Table 3: Evaluation on understanding benchmarks.
@@ 4.3 Results on Visual Understanding
LLaVA [2] 21.6 57.2 76.3 38.7 33.5 34.1
LLaVA-v1.5 [60] 27.6 63.4 85.9 64.3 58.6 35.4
InstructBLIP-7B [40] 31.2 60.9 86.1 36.0 53.4 30.6
Emu3-Chat [4] 31.7 49.2 85.2 58.5 68.2 31.6
Qwen-VL-Chat [1] 36.8 67.7 74.9 60.6 58.2 37.0
Qwen2.5-VL-7B [61] 52.9 91.0 85.9 83.2 77.0 58.6
@@
Chameleon-7B [31] 17.1 26.6 19.4 15.4 30.5 22.4
Show-o-256 [5] 73.8 25.1
Show-o-512 [5] 80.0 26.7
ILLUME [12] 88.5 65.1 72.9 38.2
TokenFlow-XL [58] 86.8 68.9 68.7 38.7
LWM [58] 75.2 -
VILA-U [30] 85.8 59.0
Janus-Pro-7B [7] 37.0 74.0 86.8 79.6 71.9 41.1
Janus-Pro-7B + DSR 38.9 81.5 86.6 80.1 71.9 41.1
4.4 Ablation Study
We conduct ablation studies to investigate the effectiveness of the proposed unified optimization strategy on the two capabilities and the effectiveness of the dual self-reward. We compare the performance of the following setups: 1) the baseline model (Janus-Pro-7B) without any optimization; 2) only optimizing the generation capability; 3) only optimizing the understanding capability; 4) jointly optimizing the two capabilities. These setups are evaluated on the T2I-CompBench, HallusionBench, and LLaVABench benchmarks. The default optimization method is SimPO [35]. The results are shown in Table 4.
Table 4: Ablation study of unified optimization manner on T2I-CompBench, HalluBench, and LLaVABench.
@@ 4.4 Ablation Study
The results show that optimizing only the generation capability significantly improves performance on the T2I-CompBench benchmark, while this setup offers limited gains on the HalluBench and LLaVABench benchmarks. Conversely, optimizing only the understanding capability yields improvements on HalluBench and LLaVABench but slightly degrades generation performance. The unified optimization approach—jointly optimizing both capabilities—achieves a balanced and substantial improvement across all benchmarks. These results demonstrate the effectiveness of the proposed unified optimization strategy in enhancing both generation and understanding capabilities without compromising one for the other.
@@
4.5 Investigation of Different Optimization Strategies
@@
We have demonstrated that the proposed dual self-reward mechanism and the unified optimization strategy are effective in jointly improving the generation and understanding capabilities in the previous sections. To further stimulate the potential of the proposed approach, we explore a different optimization strategy by training two separate models for the two capabilities and optimizing them alternately. Specifically, we train one generation model and one understanding model in an adversarial manner, where each model is trained based on the other one’s dual rewards. At each epoch, we freeze one model and update the other one, and vice versa.
Figure 3: Comparison of optimization strategies.
Figure 4: Comparison of optimization methods.
@@
The results are shown in Figure 4. We call these two models ‘separate models’ and the model trained with the unified optimization strategy ‘unified model’. It can be observed that the performance of the separate models is better than the unified model on each type of task. This indicates that the better understanding model provides better dual rewards for the generation model, and vice versa. By this the two models can be trained in an adversarial-like manner, where one focuses on generation and the other on understanding, to mutually enhance their performance with better reward signals.
@@
In addition, we also compare two different optimization methods, including the SimPO [35] and GRPO [33] in our framework on T2I-CompBench. The results can be seen in Figure 4, which shows that SimPO is slightly better than GRPO. This is likely because DSR contains a certain level of noise, while SimPO selects only the samples with the highest and lowest rewards, naturally serving as a denoising process.
Figure 7: Evolution on LLaVABench during training.
4.6 Evolution of Capabilities with Training
@@
We test the intermediate checkpoints to investigate the evolution of the capabilities of the model during training. We conduct this experiment on T2I-CompBench and LLaVABench. The results are shown in LABEL:fig:curve and Figure 7.
@@
We can observe that the generation capability improves steadily over the training process, while the understanding capability improves at the beginning and stabilizes earlier. This indicates that the understanding of the baseline model is already strong and thus is easy to converage, while the generation capability can be improved further.
5 Conclusion

### v3

@@ 4.1. Experimental Setup
We used the training set of T2I-CompBench (Huang et al., 2023) for the text-to-image generation part, which contains 5,600 text prompts. For the visual understanding part, we randomly sampled 2,800 images from the JourneyDB (Sun et al., 2023) and COCO118K (Li et al., 2024c) derived from COCO Caption (Lin et al., 2014) respectively. It is worth noting that the image and text training data are non-parallel, meaning that we did not use any annotated image-text pairs to train the model.
@@
We employ the unified LMMs with different levels of capabilities as our backbones, including Janus-Pro-7B and Janus-Pro-1B (Chen et al., 2025), optimizing them with DSR in a self-supervised manner. We compare our tuned model with state-of-the-art (SOTA) models on visual understanding and generation tasks, including models that are designed only for understanding such as LLaVA (Liu et al., 2023) and InstructBLIP-7B (Dai et al., 2023), models designed only for generation tasks like CoMat (Jiang et al., 2024) and LDM (Rombach et al., 2022), as well as unified LMMs that can perform both tasks such as Chameleon (Lu et al., 2023) and Show-o (Xie et al., 2024). All baselines can be found in Table 1, Table 2, and Table 3.
@@
We evaluate our method on visual understanding and generation benchmarks and follow their default settings. For visual generation, we use T2I-CompBench (Huang et al., 2023) and GenEval (Ghosh et al., 2023) as the evaluation benchmarks. We test on their original text prompts without any modification. T2I-CompBench comprises 6,000 compositional text prompts evaluating 3 categories (attribute binding, object relationships, and complex compositions) and 6 sub-categories (color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions). GenEval (Ghosh et al., 2023) includes 6 tasks (single object, two objects, counting, colors, position, and color attributes). For visual understanding, we use HallusionBench (denoted as HalluBench) (Zhao et al., 2023), LLaVABench, POPE (Li et al., 2023), MMB (Liu et al., 2024a), and SEEDBench-IMG (denoted as SEED) (Li et al., 2024a).
4.2. Results on Visual Generation
We compare our models with the SOTA diffusion models and unified LMMs on the T2I-CompBench and GenEval benchmarks. The results are shown in Table 1 and Table 2, respectively. Our trained models are denoted as “Janus-Pro-7B + SUDER” and “Janus-Pro-1B + SUDER”, where the models are optimized with DSR and SimPO in the unified strategy.
The results show that our models achieve substantial improvements over the baseline models. For Janus-Pro-7B, there is an average increase of 11.68% on T2I-CompBench and overall 5% on GenEval, while Janus-Pro-1B achieves an average increase of 10.5% on T2I-CompBench and 4% on GenEval. On T2I-CompBench, our tuned models remarkably improve the performance in all categories, especially in attribute binding with an average improvement of 20% and 36%, respectively, for Janus-Pro-7B and Janus-Pro-1B. In spatial relationships and complex compositions, our tuned Janus-Pro-7B also achieve impressive increases of 5% and 3% respectively, and Janus-Pro-1B achieves 17% and 10% respectively. Our tuned model demonstrates the best performance in shape binding, texture binding, and complex compositions, and achieves the second best performance in color binding. On GenEval, our models improve over the baseline model, especially in counting and color attributes with an increase of 13% and 7% respectively for Janus-Pro-7B, and in counting and and position with an increase of 6% and 12% respectively for Janus-Pro-1B. The larger one outperforms the diffusion models overall, achieving the best performance. These results demonstrate the effectiveness and superiority of our method in improving the text-to-image generation capabilities of LMMs, despite the fact that our model is optimized in a completely self-supervised manner without any external supervision or additional data.
Table 1. Evaluation on T2I-CompBench. Und. and Gen. denote “understanding” and “generation”, respectively. Higher ($\uparrow$) values indicate better performance. The best score is in bold, with the second best score underlined. Line in blue is our model.
@@ 4.2. Results on Visual Generation
StrucDiffusion (Feng et al., 2022) 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
CompDiffusion (Liu et al., 2022) 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Attend&Excite (Chefer et al., 2023) 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
PixArt-$\alpha$ (Chen et al., 2024a) 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
CoMat (Jiang et al., 2024) 0.7827 0.5329 0.6468 0.2428 0.3187 0.3680
SD-v1.5 (Rombach et al., 2022) 0.3758 0.3713 0.4186 0.1165 0.3112 0.3047
SD-XL-base-1.0 (Podell et al., 2023) 0.5879 0.4687 0.5299 0.2131 0.3119 0.3237
FLUX.1 (Labs, 2024) 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703
@@
Show-o (Xie et al., 2024) 0.56 0.41 0.46 0.20 0.30 0.29
Emu3 (Wang et al., 2024b) 0.7544 0.5706 0.7164
Janus-Pro-1B (Chen et al., 2025) 0.3542 0.2291 0.2843 0.0756 0.2809 0.2693
@@
Janus-Pro-7B (Chen et al., 2025) 0.6426 0.3487 0.4848 0.2061 0.3086 0.3510
[…]
Table 2. Evaluation on GenEval.
@@
LlamaGen (Sun et al., 2024b) 0.71 0.34 0.21 0.58 0.07 0.04 0.32
LDM (Rombach et al., 2022) 0.92 0.29 0.23 0.70 0.02 0.05 0.37
SDv1.5 (Rombach et al., 2022) 0.97 0.38 0.35 0.76 0.04 0.06 0.43
PixArt-$\alpha$ (Chen et al., 2024a) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.1 (Rombach et al., 2022) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 (Ramesh et al., 2022) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu3-Gen (Wang et al., 2024b) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL (Podell et al., 2023) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 (Betker et al., 2023) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (Esser et al., 2024) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
@@
SEED-X (Ge et al., 2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Show-o (Xie et al., 2024) 0.95 0.52 0.49 0.82 0.11 0.28 0.53
D-DiT (Li et al., 2024b) 0.97 0.80 0.54 0.76 0.32 0.50 0.65
LWM (Liu et al., 2024c) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
Transfusion (Zhou et al., 2024) 0.63
ILLUME (Wang et al., 2024a) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
TokenFlow-XL (Liu et al., 2024c) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Chameleon (Lu et al., 2023) 0.39
Janus (Wu et al., 2024b) 0.95 0.62 0.28 0.85 0.45 0.42 0.60
Janus-Pro-1B (Chen et al., 2025) 0.99 0.82 0.48 0.90 0.62 0.57 0.73
@@
Janus-Pro-7B (Chen et al., 2025) 0.97 0.88 0.57 0.90 0.77 0.64 0.79
[…]
4.3. Results on Visual Understanding
For visual understanding, we test the models on the prevalent benchmarks, including HallusionBench (denoted as HalluBench) (Zhao et al., 2023), LLaVABench, POPE (Li et al., 2023), MMB (Liu et al., 2024a), and SEEDBench-IMG (denoted as SEED) (Li et al., 2024a), which are widely used to evaluate the understanding capabilities of LMMs. The experimental results can be found in Table 3.
On HallusionBench and LLaVABench, our proposed method achieves substantial improvements over Janus-Pro-7B, with an increase of 1.9% and 7.5%, respectively; and tuned Janus-Pro-1B achieves an increase of 1.4% and 0.3%, respectively. On other benchmarks, our models keep the performance of Janus-Pro-7B, which is already a strong baseline; and slightly improve the performance of Janus-Pro-1B. The results demonstrate that our method can improve the understanding capabilities of LMMs and maintain the original advanced performance, while remarkably improving the generation capabilities.
Table 3. Evaluation on understanding benchmarks.
@@ 4.3. Results on Visual Understanding
LLaVA (Liu et al., 2023) 21.6 57.2 76.3 38.7 33.5
LLaVA-v1.5 (Liu et al., 2024b) 27.6 63.4 85.9 64.3 58.6
InstructBLIP-7B (Dai et al., 2023) 31.2 60.9 86.1 36.0 53.4
Emu3-Chat (Wang et al., 2024b) 31.7 49.2 85.2 58.5 68.2
Qwen-VL-Chat (Bai et al., 2023) 36.8 67.7 74.9 60.6 58.2
Qwen2.5-VL-7B (Bai et al., 2025) 52.9 91.0 85.9 83.2 77.0
@@
Chameleon-7B (Lu et al., 2023) 17.1 26.6 19.4 15.4 30.5
Show-o-256 (Xie et al., 2024) 73.8
Show-o-512 (Xie et al., 2024) 80.0
ILLUME (Wang et al., 2024a) 88.5 65.1 72.9
TokenFlow-XL (Liu et al., 2024c) 86.8 68.9 68.7
LWM (Liu et al., 2024c) 75.2 -
VILA-U (Wu et al., 2024d) 85.8 59.0
Janus-Pro-1B (Chen et al., 2025) 30.4 69.5 84.7 75.6 68.2
Janus-Pro-1B + SUDER 31.8 69.8 84.8 75.9 68.6
Janus-Pro-7B (Chen et al., 2025) 37.0 74.0 86.8 79.6 71.9
Janus-Pro-7B + SUDER 38.9 81.5 86.6 80.1 71.9
4.4. Ablation Study
We conduct ablation studies to investigate the effectiveness of the proposed unified optimization strategy on the two capabilities and the effectiveness of the dual self-reward. We compare the performance of the following setups: 1) the baseline model (Janus-Pro-7B) without any optimization; 2) only optimizing the generation capability; 3) only optimizing the understanding capability; 4) jointly optimizing the two capabilities. These setups are evaluated on the T2I-CompBench, HallusionBench, and LLaVABench benchmarks. The default optimization method is SimPO (Meng et al., 2024). The results are shown in Table 4.
Table 4. Ablation study of unified optimization manner on T2I-CompBench, HalluBench, and LLaVABench.
@@ 4.4. Ablation Study
The results show that optimizing only the generation capability significantly improves performance on the T2I-CompBench benchmark, while this setup offers limited gains on the HalluBench and LLaVABench benchmarks. Conversely, optimizing only the understanding capability yields improvements on HalluBench and LLaVABench but slightly degrades generation performance. The unified training approach—jointly optimizing both capabilities—achieves a balanced and substantial improvement across all benchmarks. These results demonstrate the effectiveness of the proposed unified training strategy in enhancing both generation and understanding capabilities without compromising one for the other.
@@
4.5. Investigation of Different Training Strategies
@@
We have demonstrated that the proposed dual self-reward mechanism and the unified training strategy are effective in jointly improving the generation and understanding capabilities in the previous sections. To further stimulate the potential of the proposed approach, we explore a different training strategy by training two separate models for the two capabilities and optimizing them alternately. Specifically, we train one generation model and one understanding model in an adversarial manner, where each model is trained based on the other one’s dual rewards. At each epoch, we freeze one model and update the other one, and vice versa.
Figure 3. Comparison of optimization strategies.
Figure 4. Comparison of optimization methods.
@@
The results are shown in Figure 4. We call these two models ‘separate models’ and the model trained with the unified training strategy ‘unified model’. It can be observed that the performance of the separate models is better than the unified model on each type of task. This indicates that the better understanding model provides better dual rewards for the generation model, and vice versa. By this the two models can be trained in an adversarial-like manner, where one focuses on generation and the other on understanding, to mutually enhance their performance with better reward signals.
(a) color
(b) shape
(c) texture
Figure 5. Evolution on T2I-CompBench during training.
Figure 6. Evolution on LLaVABench.
@@
In addition, we also compare two different optimization methods, including the SimPO (Meng et al., 2024) and GRPO (Shao et al., 2024) in our framework on T2I-CompBench. The results can be seen in Figure 4, which shows that SimPO is slightly better than GRPO. This is likely because DSR contains a certain level of noise, while SimPO selects only the samples with the highest and lowest rewards, naturally serving as a denoising process.
4.6. Comparison between DSR and CLIP
To further validate the effectiveness of our DSR, we compare it with CLIP (Radford et al., 2021), which is often used as the reward model in supervised training (Wang et al., 2025). We only optimize the visual generation capability of Janus-Pro-7B with SimPO, and the results are shown in Table 5.
Table 5. Comparison between DSR and CLIP.
[…]
Janus-Pro-7B 0.6426 0.3487 0.4848 37.0 74.0
Janus-Pro-7B + CLIP 0.7700 0.5609 0.7320 36.9 74.0
Janus-Pro-7B + DSR 0.7824 0.5786 0.7272 37.3 77.3
4.7. Evolution of Capabilities within Training
@@
We test the intermediate checkpoints to investigate the evolution of the capabilities of the model during training of Janus-Pro-7B. We conduct this experiment on T2I-CompBench and LLaVABench. The results are shown in Figure 6 and Figure 6.
@@
We can observe that the generation capability improves steadily over the training process, while the understanding capability improves at the beginning and stabilizes earlier. This indicates that the understanding of the baseline model is already strong and thus is easy to converge, while the generation capability can be improved further.
5. Conclusion